# On the Selection Stability of Stability Selection and Its Applications

### v2

@@ 4 Results
The source code used for the paper is accessible through the following GitHub repository: https://github.com/MahdiNouraie/Stable-Stability-Selection. Furthermore, the stabplot R package, which facilitates the use of the two plots introduced in this paper, is available through https://github.com/MahdiNouraie/stabplot.
@@ Acknowledgments
Mahdi Nouraie was supported by the Macquarie University Research Excellence Scholarship (20213605). Samuel Muller was supported by the Australian Research Council Discovery Project Grant (DP230101908).
We gratefully acknowledge Dr. Connor Smith for his guidance and support as co-supervisor of Mahdi Nouraie’s PhD research.

### v3

@@ 4 Results
The source code used for the paper is accessible through the following GitHub repository: https://github.com/MahdiNouraie/Stable-Stability-Selection. Furthermore, the stabplot R package, which facilitates the use of the methodology introduced in this paper, is available through https://github.com/MahdiNouraie/stabplot.
@@ Acknowledgments
Mahdi Nouraie was supported by the Macquarie University Research Excellence Scholarship (20213605). Samuel Muller was supported by the Australian Research Council Discovery Project Grant (DP230101908). We gratefully acknowledge Dr. Connor Smith for his guidance and support as co-supervisor of Mahdi Nouraie’s PhD research.